OpenStudy (anonymous):

rationalize the denominator 3-sqrtx/2+sqrtx

OpenStudy (ashish56):

\[[6-5\sqrt{x}+x]/4-x\]

OpenStudy (anonymous):

i dont understand that

OpenStudy (ashish56):

if u have got 2 different terms in denominator then u simply multiply both up and down by changing the sings here u can see: \[[3-\sqrt{x}]/[2-\sqrt{x}]\]=\[[(3-\sqrt{x})(2-\sqrt{x})]/[(2-\sqrt{x})(2+\sqrt{x})]\]

OpenStudy (ashish56):

now expand the brackets

OpenStudy (anonymous):

oh

jhonyy9 (jhonyy9):

this is understandably ?

OpenStudy (anonymous):

is that exactly what i type in the answer box?

jhonyy9 (jhonyy9):

fly ! when there is on denominator 2+sqrtx so than if you multiply numerator and denominator by 2-sqrtx you will get on denominator tha same like formula a2 -b2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

this is ok ?

OpenStudy (anonymous):

yes thank you

jhonyy9 (jhonyy9):

so than will get on denominator 4-x

jhonyy9 (jhonyy9):

ok good luck bye